Determine the amplitude and period of the function y=sin4x
Found 2 solutions by lwsshak3, Alan3354:
Answer by lwsshak3(11628)

Determine the amplitude and period of the function y=sin4x
Basic equation for sin function: y=Asin(Bx-C), A=amplitude, period=2π/B, phase shift=C/B
For given equation: y=sin4x
A=1
B=4
amplitude=1
Period=2π/4=π/2

Answer by Alan3354(69443)

Determine the amplitude and period of the function y=sin4x
------------
Amplitude = 1
Period = 2pi/4 = pi/2 radians